xx FleshInfo

FleshInfo

Gabrielle Allen

Date

Abstract

Demonstrates obtaining information about computational infrastructure from the flesh

1 Purpose

This thorn demonstrates using flesh interfaces to obtain information about the computational infrastructure available to a simulation. In this case, the information obtained is simply reported to standard output. This thorn doesn’t exhaust all the available information, although hopefully it will be expanded to show all the applicable APIs in practise. Notable exceptions at the moment include information about parameters, scheduled functions and grid variables.

2 Reduction Operators

FleshInfo provides a list or registered reduction operators, along with the thorn and implementation which registered them.

3 Interpolation Operators

FleshInfo provides a list or registered interpolation operators, along with the thorn and implementation which registered them.

4 IO Methods

FleshInfo provides a list or registered IO methods, along with the thorn and implementation which registered them.

5 Coordinates

FleshInfo reports on

Registered coordinate systems, their dimensions and the thorn and implementation which registered them

For each coordinate system, a list of coordinates, their directions, the computational range and the physical index range.

6 Parameters

7 Interfaces

General Implements: fleshinfo

8 Schedule This section lists all the variables which are assigned storage by thorn CactusExamples/FleshInfo. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxINITIAL x interpinfo x information about interpolation operators

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxINITIAL x reduceinfo x information about reduction operators x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxINITIAL x ioinfo x information about io methods x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxINITIAL x coordinfo x information about coordinate systems x

11 x 11 Language: 11 c
11 x 11 Type:  11 function